# Energy-Efficient Port Selection and Beamforming Design for Integrated Data and Energy Transfer Assisted by Fluid Antennas

### v1

@@ V-B Evaluation of long term performance
In this section, we evaluate the performance of C-SAC algorithm based on Python. The voltage difference $\Delta\phi$ and moving power $P_{u}$ are set as 0.1 V and 2 W, respectively. In Fig. 6, we evaluate the convergence of our proposed C-SAC algorithm, where the WET efficiency reward, achievable throughout reward, harvested energy reward as well as one FA’s normalized position are depicted in each transmission frame. The original curves represent the exact performance of each transmission frame, while the smoothed curves are obtained using Savitzky-Golay (SG) filter to show trends clearly [35]. During the first 1280 transmission frames, the agent collects experience using the actor network with initial weights. In the subsequent 12000 transmission frames, the actor network and critic networks are trained by using the sampled experience stored in the replay buffer. Observe from Fig. 6 at the beginning of 8000 transmission frames, the smoothed reward of WET efficiency and harvested energy are always lower than zero, which indicates that the long term WET efficiency constraint and harvested energy constraint are not satisfied. After the 8000-th transmission frame, all smoothed rewards vary around the zero line, which demonstrate our C-SAC algorithm converges. In addition, the normalized position of FA become more optimized as the network updates.
@@
Fig. 7 depicts the average WET efficiency versus the number of samples $L_{s}$ with different learning rate of critic networks. The average WET efficiency is calculated as the mean value over 1000 transmission frames by executing the port selection policy provided by the well-trained C-SAC agent. Observe from Fig. 7 that with the increasing number of samples, the average WET efficiency are improved, except when the learning rate of critic network is 1e-2, which indicates that inadequate sampling of FA channel may prevent the agent from capturing key positions of the FA, leading to the output of a poor port selection policy. When the number of samples reaches 6, the average WET efficiency almost approaches its optimal value. However, when the learning rate of critic network is set to 1e-2, the average WET efficiency first increases and then decreases with the number of samples. This is because a larger learning rate and sample number make the model harder to converge, resulting in a decline in average WET efficiency. In addition, when the critic networks have a learning rate of 1e-3, the average WET efficiency outperforms other learning rates, since a smaller learning rate is prone to lead the model to converge to a local optimum. Therefore, we set the learning rate of 1e-3 as the optimal learning rate of critic networks.
@@
Fig. 8 compares the average WET efficiency performance between long-term optimization and short-term optimization under different moving power $P_{u}$. Observe from Fig. 8 that the average WET efficiency of both optimization schemes decrease with the transmission throughput requirement, which reveals the trade-off between WDT performance and WET performance. In addition, compared to short-term optimization, the long-term optimization based on C-SAC algorithm achieves a better average WET efficiency. This is because that the short-term optimization focuses on achieving the best instant WET efficiency in each transmission frame, which may result in a lower WET efficiency in subsequent transmission frame. In contrast, long-term optimization leverages the C-SAC agent to capture the time-varying characteristics of the channel, enabling more proactive and forward-looking port selection decisions. Moreover, a larger moving power of FA results in a lower WET efficiency for both long-term optimization and short-term optimization, primarily due to the elevated energy consumption associated with higher FA movement.
@@ VI Conclusion
In this paper, we investigated a fluid antenna-assisted IDET system while accounting for port switching delay and energy consumption. The beamforming vector and port selection are jointly optimized to maximize the WET efficiency from the short-term and long-term perspective, subject to the constraints of achievable throughput and energy harvesting amount. In the short-term WET efficiency optimization, the suboptimal solutions are obtained by alternately optimizing the port selection and beamforming vector using FPP-SCA and SDR. In the long-term WET efficiency optimization, a novel C-SAC algorithm is proposed to find the feasible policy for the transformed constraint satisfaction problem. Simulation results demonstrate that the short-term WET efficiency of our proposed scheme outperforms that of the energy harvesting oriented scheme as well as delay-insensitive scheme. Furthermore, the long-term WET efficiency obtained by C-SAC algorithm surpasses that of short-term optimization.

### v2

@@ V-B Evaluation of long term performance
In this section, we evaluate the performance of C-SAC algorithm based on Python. The C-SAC agent interacts with a simulated time-varying wireless environment over 12,000 consecutive transmission frames, each corresponding to one channel coherence interval. The agent continuously updates its policy during these frames, capturing the dynamics of realistic online learning deployment. The voltage difference $\Delta\phi$ and moving power $P_{u}$ are set as 0.1 V and 2 W, respectively. In Fig. 6, we evaluate the convergence of our proposed C-SAC algorithm, where the WET efficiency reward, achievable throughout reward, harvested energy reward as well as one FA’s normalized position are depicted in each transmission frame. The original curves represent the exact performance of each transmission frame, while the smoothed curves are obtained using Savitzky-Golay (SG) filter to show trends clearly [43]. During the first 1280 transmission frames, the agent collects experience using the actor network with initial weights. In the subsequent 12000 transmission frames, the actor network and critic networks are trained by using the sampled experience stored in the replay buffer. Observe from Fig. 6 at the beginning of 8000 transmission frames, the smoothed reward of WET efficiency and harvested energy are always lower than zero, which indicates that the long term WET efficiency constraint and harvested energy constraint are not satisfied. After the 8000-th transmission frame, all smoothed rewards vary around the zero line, which demonstrates our C-SAC algorithm converges. In addition, the normalized position of FA becomes more optimized as the network updates.
@@
Fig. 7 depicts the average WET efficiency versus the number of samples $L_{s}$ with different learning rate of critic networks. The average WET efficiency is calculated as the mean value over 1000 transmission frames by executing the port selection policy provided by the well-trained C-SAC agent. Observe from Fig. 7 that with the increasing number of samples, the average WET efficiency is improved, except when the learning rate of critic network is 1e-2, which indicates that inadequate sampling of FA channel may prevent the agent from capturing key positions of the FA, leading to the output of a poor port selection policy. When the number of samples reaches 6, the average WET efficiency almost approaches its optimal value. However, when the learning rate of critic network is set to 1e-2, the average WET efficiency first increases and then decreases with the number of samples. This is because a larger learning rate and sample number make the model harder to converge, resulting in a decline in average WET efficiency. In addition, when the critic networks have a learning rate of 1e-3, the average WET efficiency outperforms other learning rates, since a smaller learning rate is prone to lead the model to converge to a local optimum. Therefore, we set the learning rate of 1e-3 as the optimal learning rate of critic networks.
@@
Fig. 8 compares the average WET efficiency performance between long-term optimization and short-term optimization under different moving power $P_{u}$. Observe from Fig. 8 that the average WET efficiency of both optimization schemes decrease with the transmission throughput requirement, which reveals the trade-off between WDT performance and WET performance. In addition, compared to short-term optimization, the long-term optimization based on C-SAC algorithm achieves a better average WET efficiency. This is because the short-term optimization focuses on achieving the best instant WET efficiency in each transmission frame, which may result in lower WET efficiency in subsequent transmission frames. In contrast, long-term optimization leverages the C-SAC agent to capture the time-varying characteristics of the channel, enabling more proactive and forward-looking port selection decisions. Moreover, a larger moving power of FA results in a lower WET efficiency for both long-term optimization and short-term optimization, primarily due to the elevated energy consumption associated with higher FA movement.
@@ VI Conclusion
In this paper, we investigated a fluid antenna-assisted IDET system while accounting for port switching delay and energy consumption. The beamforming vector and port selection are jointly optimized to maximize the WET efficiency from the short-term and long-term perspective, subject to the constraints of achievable throughput and energy harvesting amount. In the short-term WET efficiency optimization, the suboptimal solutions are obtained by alternately optimizing the port selection and beamforming vector using FPP-SCA and SDR. In the long-term WET efficiency optimization, a novel C-SAC algorithm is proposed to find the feasible policy for the transformed constraint satisfaction problem. Simulation results demonstrate that the short-term WET efficiency of our proposed scheme outperforms that of the energy harvesting oriented scheme as well as delay-insensitive scheme. Moreover, although the DRL training requires a relatively long convergence time, this cost is negligible compared to the typical operational lifetime of low-power wireless devices such as IoT sensors. The training overhead is acceptable and can be amortized over time. In addition, our results show that the converged DRL policy achieves significantly higher WET efficiency than short-term optimization schemes, highlighting its effectiveness for long-term energy management. This makes the proposed DRL-based solution well-suited for sustainable deployment in FAS assisted IDET system.